Grade 12Electric Current

How would you connect (parallel and series) 24 cells each of internal resistance 1 ohm so as to get the maximum output across the load of 10 ohm?

Yash
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
Dear Student,
At the point when associated in parallel, the net resistance= R 
1/R = 1/10 + 1/24 
=>R=240/34  
At the point when in arrangement net resistance = 34 
Consequently more load is the point at which the association is arrangement. 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)